OpenStudy (anonymous):

PLEASE HELP! simplify: v-1/v divided by v+1/v

OpenStudy (anonymous):

what is v ?

OpenStudy (anonymous):

if its 0 then i solve it

OpenStudy (turingtest):

^what?

OpenStudy (anonymous):

wasnt given any value for v

OpenStudy (anonymous):

(v-1/v)/(v+1/v) = (v^2 - 1)/(v^2 +1)

OpenStudy (anonymous):

multiply top and bottom by v to clear the fractions

OpenStudy (turingtest):

and if you want difference of squares too\[(v-1)(v+1)/(v^2+1)\]

OpenStudy (anonymous):

\[\frac{v-\frac{1}{v}}{v+\frac{1}{v}}\times \frac{v}{v}\] \[\frac{v^2-1}{v^2+1}\]

OpenStudy (anonymous):

Thanks guys! You are on the money!